OpenStudy (annipuppi):

Help me work out this problem. I think I'm doing something wrong. 9|9-8x|=2x+3

OpenStudy (annipuppi):

I solved for the equations 9(-9+8x)=2x+3 and 9(9-8x)=2x+3 and got x=14 and x=13 but I don't think that's right...

OpenStudy (asnaseer):

You are given:\[9|9-8x|=2x+3\]which means:\[|9-8x|=\frac{2x+3}{9}\]agreed so far?

OpenStudy (annipuppi):

yeah

OpenStudy (asnaseer):

ok, now suppose you were told that:\[|y|=7\]what does that tell you about y?

OpenStudy (annipuppi):

Umm I don't know?

OpenStudy (asnaseer):

what values could y take on to satisfy this?

OpenStudy (annipuppi):

-7 or 7

OpenStudy (asnaseer):

perfect

OpenStudy (asnaseer):

so we have:\[|9-8x|=\frac{2x+3}{9}\]which tells us:\[9-8x=\frac{2x+3}{9}\]or:\[9-8x=-\frac{2x+3}{9}\]make sense now?

OpenStudy (annipuppi):

Yeah it does, thank you

OpenStudy (asnaseer):

yw :)

OpenStudy (annipuppi):

I still think I'm doing something wrong. Will you work them out with me? @asnaseer

OpenStudy (asnaseer):

Please list your steps so that I can spot where you may have made a mistake

OpenStudy (asnaseer):

the first step would be to get rid of the fractions by multiplying both sides by 9 to get these two equations to solve:\[81-72x=2x+3\tag{1}\]\[81-72x=-2x-3\tag{2}\]

OpenStudy (annipuppi):

9-8x=2x+3/9 9-6x=3/9 -6x=-84/9 x=14/9 for the first one

OpenStudy (asnaseer):

that is not correct

OpenStudy (asnaseer):

it is NOT: 9-8x=2x+3/9 BUT this: 9-8x=(2x+3)/9

OpenStudy (annipuppi):

so what should my first step be?

OpenStudy (asnaseer):

see what I typed above ^^^

OpenStudy (asnaseer):

where I said "the first step would be to get rid of the fractions..."

OpenStudy (asnaseer):

lets just look at the first one...

OpenStudy (annipuppi):

okay, that's where I was getting confused so I'll go try to work them out now and if you're still online will you check them?

OpenStudy (asnaseer):